OpenStudy (anonymous):

how do you subtract: 7 3/5 - 4 5/6=

OpenStudy (anonymous):

well, you could do this 7 3/5 = 5*7+3 4 5/6 = 6*4+5 --------------- then solve... which is 38 - 25 = 13

OpenStudy (anonymous):

I'm showing the answer to be 2 23/30

OpenStudy (anonymous):

one sec i messed up

OpenStudy (anonymous):

yep ur correct, Its 2 23/30, sorry i solved it the wrong way

OpenStudy (anonymous):

can u show me your work, I have the answer but don't know how to show my work